# Fermionic Basis in Conformal Field Theory: The Free Fermion Point

### v1

@@ Abstract
In this work, we use the master function approach developed in [3] to describe the CFT limit of the six-vertex model at the free fermion point. Using the ODE/IM correspondence, we obtain an explicit form of the master function. This allows us to compute the asymptotic expansion of the function $\omega(\lambda,\mu)$ describing the expectation values of the fermionic basis operators [4, 5, 11]. As a result, we describe the entire Virasoro module of the corresponding CFT, including the integrals of motion as well.
@@ 1 Introduction
where $\mathcal{W}_{\alpha-s,s}$ is the subspace of quasi-local operators of the spin $s$ with the shifted $\alpha$-parameter. The operators forming the fermionic basis are defined as formal power series of the spectral parameter $\zeta^{2}$ at $1$ and have the block structure
@@
As it was discussed in [7], the functional (12) does not depend on the concrete choice of the screening operator $Y_{\mathbf{M}}^{(-s)}$. In case of the infinite lattice, one can also change the boundary conditions and, instead of taking the traces in the right-hand side of (12), insert two one-dimensional projectors $|\kappa\rangle\langle\kappa|$ and $|\kappa+\alpha-s,s\rangle\langle\kappa+\alpha-s,s|$ at the boundary, where the vector $|\kappa\rangle$ is the eigenvector of the twisted transfer matrix $T_{\mathbf{M}}(\zeta,\kappa)$ with the maximal eigenvalue $T(\zeta,\kappa)$ in the zero spin sector. Similarly, the vector $|\kappa+\alpha-s,s\rangle$ is the eigenvector of the twisted transfer matrix $T_{\mathbf{M}}(\zeta,\kappa+\alpha-s)$ with the maximal eigenvalue $T(\zeta,\kappa+\alpha-s,s)$ in the sector with spin $s$. Then
@@
The function $\omega$ will be the main object of our study. Generally speaking, its definition is quite involved, but for this work we will only need its relation (46) to the so-called master function which we will introduce it in Section 2.
@@ 2.1 Baxter TQ-relation
where $\bar{l}=l\cdot C(\nu)$, $C(\nu)=\frac{\Gamma\left(\frac{1-\nu}{2\nu}\right)}{2\sqrt{\pi}}\Gamma\left(\frac{1}{2\nu}\right)\Gamma(\nu)^{\frac{1}{\nu}}$. In the current work we consider the free fermionic point corresponding to the value $\nu=\frac{1}{2}$. It means, in particular, that the functions $a(\lambda)$ and $d(\lambda)$ have the following scaling limit
@@
As it is well-known, the ground state eigenvalue of the transfer matrix $T(\lambda)$ corresponds to all Bethe roots being positive. For the spin-zero sector the twist parameter lies in the range $1<\kappa<1$ , which brings us to the following result:
@@ 2.2 A short note on the ODE/IM correspondence
2.3 The master function in the free fermion point
@@
The master function in the free fermion point is given by a solution its definition equation (43) that can be written in the following form
@@ 3 Correlation functions in the fermionic basis
where the coefficients $\omega_{i,j}(\kappa,\kappa^{\prime},\alpha)$ are to be determined. To do so, we use the relation (46) between the master function and $\omega(\lambda,\mu;\kappa,\kappa^{\prime},\alpha)$. In the scaling limit the operator $H_{\lambda}$ simplifies. Indeed, from [2] one concludes that for $\lambda^{2}\rightarrow\infty$ the function $\mathfrak{a}^{\mathrm{sc}}(\lambda,\kappa)$ rapidly decays in the upper half plane and rapidly grows in the lower half plane. It means that the second term in (47) does not contribute unless $f(\lambda)$ grows exponentially. Therefore, the relation (46) for $\omega^{\mathrm{sc}}(\lambda,\mu;\kappa,\kappa^{\prime},\alpha)$ becomes
@@
The relation (66) is of the most importance for us. Using the explicit form (125) of the master function and calculating the asymptotic of the left-hand side of (66) we can compute the coefficients $\omega_{i,j}(\kappa,\kappa^{\prime},\alpha)$ in the free fermion point.
@@ 3.1 The relation between the Virasoro and fermionic generators
One of the point of interest for us is to establish the explicit relation between the Fermionic basis and the Virasoro generators. Modulo integrals of motion it was done in [7, 9] up to level 8. In this section we show how to incorporate the integrals of motion in the free fermion case.
@@
Using the Ward-Takahashi identities to calculate the expectation values of even Virasoro generators descendents on the left-hand side of (80) once again, we come to the following result for the coefficients $y_{1,3}(\alpha)$:
@@
The coefficients $y_{3,1}$ in front of the integrals of motions in the expansion of the Fermionic basis generators $\beta_{3}\gamma_{1}$ via the Virasoro generators can be obtained from the above coefficients $y_{1,3}$ replacing $\alpha$ by $2-\alpha$:
@@
In order to proceed with higher levels, however, it is not enough to consider the three-point function with the primary fields at the boundary only. We also have to consider the descendents. It was done in [9, 1] for the case $\kappa=\kappa^{\prime}$. We hope to incorporate the particle-hole excitations in a future work.
@@ 4.1 The case of equal boundary fields
The case of equal boundary fields corresponding to the choice $\kappa=\kappa^{\prime}$ was considered before in [7] for $\frac{1}{2}\leq\nu\leq 1$. It was done with the help of Wiener–Hopf factorisation technique. Since it is not applicable to the case $\kappa\neq\kappa^{\prime}$, we do not discuss it in this work, and instead just present the final result of [7]. The asymptotic expansion of the function $\omega$ at $\lambda,\mu\rightarrow\infty$ is given by
@@ 4.2 The reflection relations
The right hand side of the above identity is invariant under $\sigma_{1}$. The first term of the left-hand side is invariant under $\sigma_{2}$, but the second one is not. The idea is to correct this adding $x(\alpha)\operatorname{\mathbf{i}}_{1}^{2}|a;0\rangle$ in such a way that it does not spoil the $\sigma_{1}$-invariance of the right-hand side. It means that the coefficient $x(\alpha)$ should satisfy the following set of equations  [^1] We take the definition of $D_{2n-1}(\alpha,\nu)$ from [7]. This definition differs from that used in [8] by a factor of $\operatorname{i}$; as a result, the l.h.s. of (96b) appears with the opposite sign compared to [8]. :
@@
Now let us discuss level 4. The construction of [8] becomes cumbersome rather quickly. However, it simplifies greatly in the free fermion point, and the integrals involved in the calculation of the coefficients in front of the integrals of motion can be calculated explicitly. We mention here some typos made in Appendix of [8]: some of the coefficients $X_{1,3}$ and $X_{3,1}$ are given with wrong overall sign. Namely, one has to perform the following change:
@@ 5 Conclusions and discussion
Let us make a small remark related to the last question. In the case of equal boundary condition, the correspondence between fermionic basis operators and Virasoro generators was established in [7, 9] modulo integrals of motion up to level 8. Let us denote the combinations of even Virasoro generators as $V_{i,j}(\alpha)$. For example, from (78) and (80), we get
@@ Acknowledgements
The authors would like thank F. Smirnov for collaboration at the early stages of this work, which originated from intensive discussions with him. HB thanks F. Göhmann for his interest and valuable discussions. SA thanks M. Minin for numerous stimulating discussions.

### v2

@@ Abstract
In this work, we use the master function approach developed in [3] to describe the CFT limit of the six-vertex model at the free fermion point. Using the ODE/IM correspondence, we obtain an explicit form of the master function. This allows us to compute the asymptotic expansion of the function $\omega(\lambda,\mu)$ describing the expectation values of the fermionic basis operators [4, 5, 11]. As a result, we describe the entire Virasoro module of the corresponding CFT, including the integrals of motion.
@@ 1 Introduction
where $\mathcal{W}_{\alpha-s,s}$ is the subspace of quasi-local operators of the spin $s$ with the shifted $\alpha$-parameter. The operators forming the fermionic basis are defined as formal power series of the spectral parameter $\zeta^{2}$ around $\zeta^{2}=1$ and have the block structure
@@
As it was discussed in [7], the functional (12) does not depend on the concrete choice of the screening operator $Y_{\mathbf{M}}^{(-s)}$. In the case of the infinite lattice, one can also change the boundary conditions and, instead of taking the traces in the right-hand side of (12), insert two one-dimensional projectors $|\kappa\rangle\langle\kappa|$ and $|\kappa+\alpha-s,s\rangle\langle\kappa+\alpha-s,s|$ at the boundary, where the vector $|\kappa\rangle$ is the eigenvector of the twisted transfer matrix $T_{\mathbf{M}}(\zeta,\kappa)$ with the maximal eigenvalue $T(\zeta,\kappa)$ in the zero spin sector. Similarly, the vector $|\kappa+\alpha-s,s\rangle$ is the eigenvector of the twisted transfer matrix $T_{\mathbf{M}}(\zeta,\kappa+\alpha-s)$ with the maximal eigenvalue $T(\zeta,\kappa+\alpha-s,s)$ in the sector with spin $s$. Then
@@
The function $\omega$ will be the main object of our study. Generally speaking, its definition is quite involved, but for this work we will only need its relation (46) to the so-called master function which we will introduce in Section 2.
@@ 2.1 Baxter TQ-relation
where $\bar{l}=l\cdot C(\nu)$, $C(\nu)=\frac{\Gamma\left(\frac{1-\nu}{2\nu}\right)}{2\sqrt{\pi}}\Gamma\left(\frac{1}{2\nu}\right)\Gamma(\nu)^{\frac{1}{\nu}}$. In the current work we consider the free fermionic point corresponding to the value $\nu=\frac{1}{2}$. This means, in particular, that the functions $a(\lambda)$ and $d(\lambda)$ have the following scaling limit
@@
As is well-known, the ground state eigenvalue of the transfer matrix $T(\lambda)$ corresponds to all Bethe roots being positive. For the spin-zero sector the twist parameter lies in the range $1<\kappa<1$ , which brings us to the following result:
@@ 2.2 A short note on the ODE/IM correspondence
2.3 The master function at the free fermion point
@@
The master function at the free fermion point is given by a solution of its defining equation (43) that can be written in the following form
@@ 3 Correlation functions in the fermionic basis
where the coefficients $\omega_{i,j}(\kappa,\kappa^{\prime},\alpha)$ are to be determined. To do so, we use the relation (46) between the master function and $\omega(\lambda,\mu;\kappa,\kappa^{\prime},\alpha)$. In the scaling limit the operator $H_{\lambda}$ simplifies. Indeed, from [2] one concludes that for $\lambda^{2}\rightarrow\infty$ the function $\mathfrak{a}^{\mathrm{sc}}(\lambda,\kappa)$ rapidly decays in the upper half plane and rapidly grows in the lower half plane. This means that the second term in (47) does not contribute unless $f(\lambda)$ grows exponentially. Therefore, the relation (46) for $\omega^{\mathrm{sc}}(\lambda,\mu;\kappa,\kappa^{\prime},\alpha)$ becomes
@@
The relation (66) is of utmost importance for us. Using the explicit form (125) of the master function, and calculating the asymptotic of the left-hand side of (66) we can compute the coefficients $\omega_{i,j}(\kappa,\kappa^{\prime},\alpha)$ at the free fermion point.
@@ 3.1 The relation between the Virasoro and fermionic generators
One of the points of interest for us is to establish the explicit relation between the Fermionic basis and the Virasoro generators. Modulo integrals of motion it was done in [7, 9] up to level 8. In this section we show how to incorporate the integrals of motion in the free fermion case.
@@
Using the Ward-Takahashi identities to calculate the expectation values of even Virasoro generators descendants on the left-hand side of (80) once again, we come to the following result for the coefficients $y_{1,3}(\alpha)$:
@@
The coefficients $y_{3,1}$ in front of the integrals of motion in the expansion of the Fermionic basis generators $\beta_{3}\gamma_{1}$ via the Virasoro generators can be obtained from the above coefficients $y_{1,3}$ replacing $\alpha$ by $2-\alpha$:
@@
In order to proceed with higher levels, however, it is not enough to consider the three-point function with the primary fields at the boundary only. We also have to consider the descendants. It was done in [9, 1] for the case $\kappa=\kappa^{\prime}$. We hope to incorporate the particle-hole excitations in a future work.
@@ 4.1 The case of equal boundary fields
The case of equal boundary fields corresponding to the choice $\kappa=\kappa^{\prime}$ was considered before in [7] for $\frac{1}{2}\leq\nu\leq 1$. It was done with the help of the Wiener–Hopf factorisation technique. Since it is not applicable to the case $\kappa\neq\kappa^{\prime}$, we do not discuss it in this work, and instead just present the final result of [7]. The asymptotic expansion of the function $\omega$ at $\lambda,\mu\rightarrow\infty$ is given by
@@ 4.2 The reflection relations
The right hand side of the above identity is invariant under $\sigma_{1}$. The first term of the left-hand side is invariant under $\sigma_{2}$, but the second one is not. The idea is to correct this adding $x(\alpha)\operatorname{\mathbf{i}}_{1}^{2}|a;0\rangle$ in such a way that it does not spoil the $\sigma_{1}$-invariance of the right-hand side. This means that the coefficient $x(\alpha)$ should satisfy the following set of equations  [^1] We take the definition of $D_{2n-1}(\alpha,\nu)$ from [7]. This definition differs from that used in [8] by a factor of $\operatorname{i}$; as a result, the l.h.s. of (96b) appears with the opposite sign compared to [8]. :
@@
Now let us discuss level 4. The construction of [8] becomes cumbersome rather quickly. However, it simplifies greatly at the free fermion point, and the integrals involved in the calculation of the coefficients in front of the integrals of motion can be calculated explicitly. We mention here some typos made in the Appendix of [8]: some of the coefficients $X_{1,3}$ and $X_{3,1}$ are given with wrong overall sign. Namely, one has to perform the following change:
@@ 5 Conclusions and discussion
Let us make a small remark related to the last question. In the case of equal boundary conditions, the correspondence between fermionic basis operators and Virasoro generators was established in [7, 9] modulo integrals of motion up to level 8. Let us denote the combinations of even Virasoro generators as $V_{i,j}(\alpha)$. For example, from (78) and (80), we get
@@ Acknowledgements
The authors would like to thank F. Smirnov for collaboration at the early stages of this work, which originated from intensive discussions with him. HB thanks F. Göhmann for his interest and valuable discussions. SA thanks M. Minin for numerous stimulating discussions.